# Optimal Hold-Out Size in Cross-Validation

### v2

@@ Abstract
Cross-validation (CV) is a standard technique used across science to test how well a model predicts new data. Data are split into $K$ “folds,” where one fold (i.e., hold-out set) is used to evaluate a model’s predictive ability, with folds cycled in standard $K$-fold CV. Researchers typically rely on conventions when choosing the hold-out size, commonly $80/20$ split, or $K=5$, even though this choice can affect inference and model evaluation. Principally, this split should be determined by balancing the predictive accuracy (bias) and the uncertainty of this accuracy (variance), which forms a tradeoff based on the size of the hold-out set. More training data means more accurate models, but fewer testing data lead to uncertain evaluation, and vice versa. The challenge is that this evaluation uncertainty cannot be identified, without strong assumptions, directly from data. We propose a procedure to determine the optimal hold-out size by deriving a finite-sample exact expression and upper bound on the evaluation uncertainty, depending on the error assumption, and adopting a utility-based approach to make this tradeoff explicit. Analyses of real-world datasets using linear regression and random forest demonstrate this procedure in practice, providing insight into implicit assumptions, robustness, and model performance. Critically, the results show that the optimal hold-out size depends on both the data and the model, and that conventional choices implicitly make assumptions about the fundamental characteristics of the data. Our framework makes these assumptions explicit and provides a principled, transparent way to select this split based on the data and model rather than convention. By replacing a one-size-fits-all choice with context-specific reasoning, it enables more reliable comparisons of predictive performance across scientific domains.
@@ Introduction
In typical data analysis, the researcher has a plethora of models or hyperparameters to choose from. For example, a researcher often has access to a set of models that generate drastically different inferences or predictions, and the researcher must choose one of these models to make informed decisions. Which model to use– i.e., model uncertainty– is a central issue in statistics and machine learning. To assess model performance, a common technique used is cross-validation (CV), where a certain portion of the data is randomly allocated to train the model, and the remaining portion is used to evaluate the model through its predictions. Cross-validation, developed simultaneously and independently by Seymour Geisser and Mervyn Stone (christensen2007conversation), is popular, and used across a broad breadth of fields (e.g., machine learning, econometrics, genomics, and epidemiology, to name just a few), because the concept and implementation are simple, while not being constrained to a specific data type or a domain, and theoretically known to produce unbiased estimates of the prediction loss (see, e.g., efron1983estimating; shao1993linear; bates2024cross). In simple terms, cross-validation tests how well a model trained on part of the data predicts the rest. Repeating this process multiple times provides an estimate of how the model will generalize to unseen data. The most common version is $K$-fold cross-validation, where the data are split into $K$ equal-sized folds, and the test set is constructed by removing one of them. In practice, $K$ is determined by convention (e.g., 80/20 split, or 5-fold CV), and does not typically differ across domain, dataset, or model. Even though this choice of $K$ directly affects selection and inference, there is no principled way to determine this split with theoretical justification. Consider a concrete example: a researcher analyzing a dataset might find that Model A outperforms Model B using 5-fold CV, but Model B beats Model A using 10-fold CV. This could occur because Model B is more complex than Model A, requiring more data to produce good forecasts, but it could also occur because, particularly if the total sample size is small, the test data happened to favor Model B. In either case, the researcher could select the wrong model and reach incorrect scientific conclusions if the split is poorly chosen. More importantly, a researcher faced with different results based on the number of folds does not have a principled way to discern which to use. Since the variability of cross-validation depends directly on the size of the hold-out set, $m$ (with $K$-fold CV corresponding to $m=N/K$), our focus is on determining the optimal hold-out size.
The difficulty in determining the hold-out size is due to the inherent tradeoff that exists in cross-validation. The more training data one has, the better the model estimation will be. Conversely, the more test data one has, the better the performance assessment of the model will be. Given a fixed total sample size, a tradeoff exists between the accuracy of the model and the uncertainty of the model evaluation. Consider an extreme case where all of the data are allocated to the training set. The model is the most accurate, using all the data to train, but there is no model evaluation (due to not having any test samples), and the uncertainty regarding the assessment is the largest. At the other extreme, if all data are allocated to the test set, the model cannot learn from the data and can only predict based on prior assumptions. The evaluation uncertainty, however, is the smallest because all of the data are used to evaluate. This tradeoff suggests that an optimal balance exists between having sufficient training data for learning and enough test data for reliable evaluation. This simple thought experiment illustrates why a universal choice of hold-out size cannot exist, since the optimal split depends on how much uncertainty one can tolerate in evaluating model accuracy.
Our goal is to formalize this tradeoff to enable a systematic choice of the optimal hold-out size. More precisely, this tradeoff is between the prediction loss of the model and its uncertainty regarding the prediction loss; similar to the bias-variance tradeoff regarding model complexity. While this tradeoff has been intuited, it has never been rigorously explicated due to the impossibility of a universal unbiased estimator for the latter uncertainty in $K$-fold CV (bengio2004no), except for specific problem settings (e.g., arlot2016choice), and only asymptotic results have been derived for the purpose of testing model performance (austern2020asymptotics; bayle2020cross; li2023asymptotics). While these results provide valuable asymptotic insights, they do not yield a usable criterion for choosing the split in a given dataset with finite samples. For our theoretical contribution, we derive a finite sample exact bound (for symmetric error) and upper bound (for more general error) to this uncertainty, rigorously linking it to the prediction loss. Given this uncertainty bound, the optimal test set size can be derived by balancing this tradeoff. We express this tradeoff using a mean-variance utility (a combined score of accuracy and reliability) of the von Neumann-Morgenstern utility function. The optimal test set size maximizes this utility function, given the researcher’s belief in the irreducible error in the data-generating process, which acts as a tuning parameter.
The theoretical analysis in this paper concerns the evaluation loss on a single hold-out set of size $m$, which represents the fundamental uncertainty of assessing a model on $m$ new observations. In practice, this choice of $m$ is implemented through standard $K$-fold cross-validation by setting $m=N/K$. $K$-fold cross-validation averages several such hold-out evaluations, reducing variance but not altering the bias-variance tradeoff that determines the optimal hold-out size.
A key insight of our approach is recognizing that any choice of hold-out size implicitly assumes the level of irreducible noise in the data-generating process. Thus, choosing 80/20 assumes one noise level and 90/10 assumes another. Current practice makes these choices by convention without examining what they imply about the fundamental characteristics of the data. We make these implicit assumptions explicit, transforming convention into a transparent, data-informed decision.
In the following, we formalize this tradeoff, derive theoretical results for its components, and apply it to real data.
@@
To demonstrate our approach in practice, we apply it to a regression problem using the Abalone dataset (abalone_1). The Abalone dataset has $N=4,177$ observations with eight covariates, and the goal is to predict the age of the abalone based on these features. We consider two models, a linear regression model and a random forest model, for comparison to estimate the optimal test set size, due to the latter being a more complex (nonlinear) model compared to the former.
@@
Abalone: Age of abalone abalone_1 Biology Rings 4,177 8
Concrete: Concrete compressive strength concrete_compressive_strength_165 Materials Engineering Concrete strength 1,030 8
Servo: Simulation of a servo system servo_87 Engineering Rise time 167 4
Diabetes: Diabetes study efron2004least Medicine Disease progression score 442 10
Income: Census income adult_2 Social Science log(Income) 32,561 5
@@ Discussion
The cross-validation procedure can be viewed as a balance between two competing goals: model accuracy and evaluation certainty. Based on this tradeoff, we devise a strategy based on multi-objective optimization to determine the optimal hold-out size in cross-validation. Our proposed strategy is easily implemented through basic cross-validation estimates and is optimized based on the researcher’s belief in the irreducible noise. Through regression exercises, we demonstrate how the optimal test set size can be estimated, and how that informs each model’s predictive ability, sensitivity to assumptions regarding the irreducible error, and how conventional splits can mislead evaluation. Our proposed method can be applied to many statistical problems, but is particularly useful (due to a tighter bound) for regression problems where the errors are symmetrically distributed and homoskedastic.
@@
Our approach has several limitations. The most important thing is the computational cost of estimating the cross-validation loss. While we have discussed reducing this cost by reducing the number of test sets in cross-validation, another approach is to use the convergence rate of each model, though this requires knowledge of this rate, which can be difficult to obtain. As discussed above, whether one reduces the number of test sets, uses the convergence rate, or SURE to avoid LOOCV, depends on the availability of computational resources and theoretical understanding.
Finally, as seen in the results of our analysis, it is not always necessary to obtain exact optimal test set sizes. Our results show that, for the Abalone data, there is a reasonable range for which little utility is lost by deviating from the optimal size. The key result is that the optimal size will rarely be LOOCV, or something like a 2-fold CV, but somewhere in between. Instead of relying on a single conventional across models and datasets, we recommend identifying a reasonable range of hold-out sizes tailored to specific problem types or data domains. This is similar to having different significance levels for statistical testing for different domains. The main takeaway is that there is no one-size-fits-all for cross-validation, and it should be considered on a domain-by-domain, dataset-by-dataset, and model-by-model basis.
@@
Our theoretical results provide an exact expression and upper bound for the evaluation uncertainty, depending on the error assumption, at a given hold-out size $m$, which is directly applicable to single-split validation and remains conservative when implemented through $K$-fold averaging. A natural direction for future work is to derive bounds that directly target the variance of the $K$-fold averaged estimator. Such results would require explicit control of cross-fold dependence, since overlapping training sets induce nonzero covariances between fold losses. One promising approach is to express $\mathrm{Var}(K\text{-CV})$ in terms of per-fold variances and average cross-fold correlations, yielding an “effective number of folds” that interpolates between the independent-fold regime (variance reduction $\approx 1/K$) and the fully dependent regime (no reduction). Establishing sharp, assumption-transparent bounds of this kind—potentially using algorithmic stability or related dependence controls—would further tighten uncertainty quantification while preserving the interpretability of our hold-out size framework.
@@
Our variance bound is derived for the evaluation loss on a single hold-out set of size $m$, and is therefore the appropriate conservative choice when a practitioner uses a single train/test split or wishes to retain a worst-case guarantee at a given hold-out size. In many applications, however, researchers report the $K$-fold averaged cross-validation estimate, which aggregates $K$ fold losses and typically has substantially smaller variance than a single hold-out evaluation. In this setting, the single-split bound remains valid but can be conservative. For practitioners who wish to approximate the uncertainty of the reported $K$-fold average more closely, we consider a simple CLT-style rescaling of the bound as a heuristic “practical” variance proxy, and evaluate its accuracy empirically. Across all tested noise regimes and hold-out sizes, this practical proxy closely tracks the Monte Carlo variance of the $K$-fold estimator well for large folds ($K\geq 4$) (Appendix Figure 9). While this heuristic does not carry the same finite-sample guarantee as the theoretical bound, it provides a useful operational approximation.
@@ A.5 Proof of Theorem 1
(Theorem 1 of Whittle_60) Let the random variables $z_{1},z_{2},\ldots,z_{n}$ be statistically independent, and each assume the values $\pm 1$ with respective probabilities $\frac{1}{2}$ . Then, for any set of real coefficients $b_{j}$, $s\geq 2$,

### v3

@@ Abstract
Cross-validation (CV) is routinely used across the sciences to select models and tune parameters, and the resulting choices are often interpreted as substantive scientific conclusions (e.g., which variables, mechanisms, or risk factors are “supported by the data”). A key part of the CV procedure– the hold-out size, or equivalently the fold count KK– is typically set by convention (e.g., 80/20, $K=5$) rather than by a principled criterion. Central to the issue is the tradeoff between training and testing: increasing the training sample size improves model accuracy, while sacrificing certainty around the accuracy itself. We formalize the tradeoff by targeting predictive performance and explicitly penalizing evaluation uncertainty, which cannot be identified from the data without additional assumptions. We derive finite-sample expressions of this evaluation uncertainty under symmetric errors and general upper bounds under broader error conditions, yielding a transparent utility-based rule for selecting the hold-out size as a function of an irreducible-noise parameter. Empirical analyses with linear regression and random forests across multiple domains, and a high-dimensional genomics application, show that (i) the choice of $K$ is dependent on the data and model. (ii) the optimal $K$ varies based on the assumption on the irreducible error, and (iii) the implied inferential conclusions can change materially as the irreducible error, and thus $K$, varies. The resulting framework replaces a one-size-fits-all convention with a context-specific, assumption-explicit choice of $K$, enabling more reliable model comparisons and downstream scientific inference.
@@ Introduction
Cross-validation (CV) is widely treated as a “neutral” way to evaluate predictive performance and choose among models or tuning parameters. In practice, however, CV is often used as a gatekeeper for scientific claims: a model choice made by CV can determine which covariates are reported as significant, which mechanisms are described as active, or which policies are recommended. In these settings, the CV procedure is not merely a computational device; it is part of the inferential pipeline. A central and rarely justified component of that pipeline is the hold-out size, or equivalently, the number of folds $K$. Most applied work defaults to convention (80/20, $K=5$, $K=10$), even though changing $K$ can change the selected model and the conclusions drawn from it.
The reason is simple: $K$ controls two competing effects. Larger training sets can yield better-fitted models (lower predictive bias), but smaller test sets yield noisier performance estimates (higher evaluation variance). When predictive performance is strongly separated across candidate procedures, these differences may not matter. However, in many modern problems– especially with flexible models, many tuning parameters, or high-dimensional sparsity– the performance curve can be nearly flat. In that regime, the choice of $K$ can act primarily on stability and selection rather than on average fit: two choices of $K$ may have essentially identical estimated loss, yet produce different selected models and, therefore, different downstream narratives. A researcher faced with conflicting CV outcomes across fold counts has no standard, principled way to decide which $K$ is appropriate.
This paper treats the choice of hold-out size as a decision problem. Our target is pure predictive performance– performance net of irreducible noise– rather than raw empirical loss, and we make explicit that any decision based on CV must confront uncertainty in the evaluation itself. The main obstacle is that evaluation uncertainty in $K$-fold CV is not universally estimable from the observed data without further assumptions (Bengio and Grandvalet, 2004). Existing theory provides valuable asymptotic descriptions of CV-based comparisons and tests (e.g., Austern and Zhou, 2025; Bayle et al., 2020; Li, 2023), and guidance for particular settings (e.g., Arlot and Lerasle, 2016), but does not yield a general finite-sample, dataset-specific criterion that transparently trades off accuracy against evaluation uncertainty.
Our contribution is to provide such a criterion by deriving finite-sample characterizations of the evaluation uncertainty (exact under symmetric errors; upper bounds more generally), linking this uncertainty to the loss, and embedding the tradeoff in an explicit mean–variance utility. This yields a transparent rule that maps an irreducible-noise parameter (reflecting the analyst’s tolerance for evaluation uncertainty) to an optimal hold-out size. In contrast to a single “recommended” $K$, the output is a robustness map: which $K$ is optimal under which noise regimes, and how conclusions change across that map.
A final practical point concerns implementation. The theoretical analysis concerns evaluation on a single hold-out set of size $m$, representing the fundamental uncertainty of assessing a fitted model on $m$ new observations. In standard $K$-fold CV, $m=N/K$ and the reported loss averages across folds; this averaging reduces variance but does not remove the underlying tradeoff induced by the hold-out size. Thus, choosing $K$ remains a scientific decision: it determines how aggressively one prioritizes training accuracy versus reliable evaluation.
The remainder of the paper formalizes this tradeoff, derives finite-sample results that make its components explicit, and demonstrates the implications on real data. In particular, we show that the optimal hold-out size depends on both the model class and the dataset, and that conventional choices implicitly encode assumptions about irreducible noise and acceptable evaluation uncertainty. By making these assumptions explicit, the proposed framework replaces convention with a principled, transparent choice of $K$ and clarifies how that choice propagates to downstream inference.
@@
To demonstrate our approach in practice, we apply it to a regression problem using the Abalone dataset (Nash et al., 1994). The Abalone dataset has $N=4,177$ observations with eight covariates, and the goal is to predict the age of the abalone based on these features. We consider two models, a linear regression model and a random forest model, for comparison to estimate the optimal test set size, due to the latter being a more complex (nonlinear) model compared to the former.
@@
Abalone: Age of abalone Nash et al. (1994) Biology Rings 4,177 8
Concrete: Concrete compressive strength Yeh (1998) Materials Engineering Concrete strength 1,030 8
Servo: Simulation of a servo system Ulrich (1986) Engineering Rise time 167 4
Diabetes: Diabetes study Efron et al. (2004) Medicine Disease progression score 442 10
Income: Census income Becker and Kohavi (1996) Social Science log(Income) 32,561 5
@@
Why $K$ matters scientifically
Mutational-signature attribution turns high-dimensional somatic mutation data into a sparse set of active mutational processes, typically represented in the 96 trinucleotide-context basis (Alexandrov et al., 2013). Because the scientific conclusions are expressed in terms of which signatures are deemed present and how strongly they contribute, any tuning choice that changes sparsity can change downstream mechanistic claims.
We illustrate this using TCGA/MC3 somatic mutation calls (Ellrott et al., 2018), converted to SBS96 spectra using standard MAF tooling (Mayakonda et al., 2018) and refit against the COSMIC reference catalogue (Sondka et al., 2024). Signature weights are estimated via $\ell_{1}$-regularized (nonnegative) regression, as in signature “refitting” approaches designed to yield sparse, interpretable assignments (Li et al., 2020; Friedman et al., 2010). We then apply our proposed utility framework: for each fold count $K$, compute the hold-out loss and select the $K$ that minimizes the corresponding risk-adjusted estimate of pure loss as a function of $\sigma^{2}$.
The results are reported in Table 4. First, the fit can be nearly insensitive to $K$: across a wide range of fold counts, the empirical hold-out loss changes only marginally. Second, despite that near-flat fit curve, the optimal $K$ can vary sharply with $\sigma^{2}$ under the pure-loss criterion, producing a step-function frontier $K$ in which larger uncertainty favors smaller $K$. This is exactly the regime in which downstream selection is most fragile. When multiple $K$ values have essentially identical predictive performance, the fold design becomes a lever on sparsity rather than accuracy.
In our TCGA analysis, this translates into different substantive conclusions about which signatures are “reliably present.” Using a stability threshold (a signature is declared present only if selected in at least 70% of tumors), the high-uncertainty regime (where the pure-loss criterion selects small $K$) excludes signatures that clear the same threshold under lower uncertainty (where the criterion selects a larger $K$). For example, SBS24, SBS39, and SBS49 cross the stability threshold when moving from the high-uncertainty optimum ($K=2$) to the lower-uncertainty optima ($K\geq 5$). Thus, varying $K$ changes not only numerical weights but the discrete set of inferred processes and, consequently, the biological narratives supported by the data.
Table 4: Optimal $K$ as a function of $\sigma^{2}$ (TCGA MC3).
$K$ $\sigma^{2}$ range (low) $\sigma^{2}$ range (high)
100 0 $7.41\times 10^{-4}$
20 $7.41\times 10^{-4}$ $6.84\times 10^{-3}$
10 $6.84\times 10^{-3}$ $4.03\times 10^{-2}$
5 $4.03\times 10^{-2}$ $3.52\times 10^{-1}$
2 $3.52\times 10^{-1}$
@@ Discussion
The cross-validation procedure can be viewed as a balance between two competing goals: model accuracy and evaluation certainty. Based on this tradeoff, we devise a strategy based on multi-objective optimization to determine the optimal hold-out size in cross-validation. Our proposed strategy is easily implemented through basic cross-validation estimates and is optimized based on the researcher’s belief in the irreducible noise. Through regression exercises, we demonstrate how the optimal test set size can be estimated, and how that informs each model’s predictive ability, sensitivity to assumptions regarding the irreducible error, and how conventional splits can mislead evaluation. Our proposed method can be applied to many statistical problems, but is particularly useful (due to an exact identity) for regression problems where the errors are symmetrically distributed and homoskedastic.
@@
Our approach has several limitations. The most important thing is the computational cost of estimating the cross-validation loss. While we have discussed ways to reduce this cost, another approach is to use the convergence rate of each model, though this requires knowledge of this rate, which can be difficult to obtain. As discussed above, whether one reduces the number of test sets, uses the convergence rate, or SURE to avoid LOOCV, depends on the availability of computational resources and theoretical understanding.
Finally, as seen in the results of our analysis, it is not always necessary to obtain exact optimal test set sizes. Our results show that, for the Abalone data, there is a reasonable range for which little utility is lost by deviating from the optimal size. The key result is that the optimal size will rarely be LOOCV, or something like a 2-fold CV, but somewhere in between. Instead of relying on a single conventional choice across models and datasets, we recommend identifying a reasonable range of hold-out sizes tailored to specific problem types or data domains. This is similar to having different significance levels for statistical testing for different domains. The main takeaway is that there is no one-size-fits-all for cross-validation, and it should be considered on a domain-by-domain, dataset-by-dataset, and model-by-model basis.
@@
Our theoretical results provide an exact expression and upper bound for the evaluation uncertainty, depending on the error assumption, at a given hold-out size $m$, which is directly applicable to single-split validation and remains conservative when implemented through $K$-fold averaging. A natural direction for future work is to derive bounds that directly target the variance of the $K$-fold averaged estimator. Such results would require explicit control of cross-fold dependence, since overlapping training sets induce nonzero covariances between fold losses. One promising approach is to express $\mathrm{Var}(K\text{-CV})$ in terms of per-fold variances and average cross-fold correlations, yielding an “effective number of folds” that interpolates between the independent-fold regime (variance reduction $\approx 1/K$) and the fully dependent regime (no reduction). Establishing sharp, assumption-transparent bounds of this kind– potentially using algorithmic stability or related dependence controls– would further tighten uncertainty quantification while preserving the interpretability of our hold-out size framework.
@@
Practical recommendations.
Based on our theoretical and empirical results, we offer the following guidance for applied researchers:
  1. 1.
    Low-noise settings (controlled experiments, high-precision measurements): The evaluation-uncertainty term is typically small, so it is often reasonable to use a larger $K=N/m$ (e.g., 10–20 folds), prioritizing training data to reduce bias.
  2. 2.
    High-noise settings (surveys, behavioral data, many social-science applications): Evaluation uncertainty can dominate, so smaller $K$ (e.g., 2–5 folds) can be preferable to stabilize model evaluation, even at the cost of reduced training size.
  3. 3.
    Complex models (random forests, neural networks, ensembles): These models can be data-hungry, so a larger $K$ (smaller hold-out) is often beneficial when noise is moderate; however, under high noise, the optimal $K$ can shift downward because uncertainty in evaluation dominates.
  4. 4.
    Simple models (linear/logistic regression): These models typically learn effectively from smaller training sets, so the optimal $K$ is often more sensitive to the assumed noise level and the desired precision of model evaluation.
  5. 5.
    When uncertain: Compute the Pareto frontier across a plausible range of $\sigma^{2}$. If the recommended $K$ is stable across $\sigma^{2}$ (e.g., K=5K=51010 over $\sigma^{2}\in[0.1,1]$), report this robustness range. If it varies substantially, report results under multiple splits and interpret conclusions as sensitive to noise assumptions.
  6. 6.
    For model comparison: Use the same $\sigma^{2}$ assumption when comparing models on the same dataset, but allow that the optimal $K$ may differ across models. When conclusions change across $K$, the Pareto frontier helps diagnose whether the disagreement reflects genuine performance differences or sensitivity to evaluation uncertainty.
As a diagnostic tool, reverse-engineering the implied $\sigma^{2}$ under conventional splits (Table 3) can reveal whether standard practice aligns with domain knowledge about data quality. Because $\sigma^{2}$ is scale-dependent, this diagnostic is most interpretable after standardizing the outcome (or otherwise fixing its measurement scale). Implausible implied values under that scale (e.g., unusually large noise for a well-controlled setting) suggest that conventional $K$ is inappropriate for the problem at hand.
@@
Our variance bound is derived for the evaluation loss on a single hold-out set of size $m$, and is therefore the appropriate conservative choice when a practitioner uses a single train/test split or wishes to retain a worst-case guarantee at a given hold-out size. In many applications, however, researchers report the $K$-fold averaged cross-validation estimate, which aggregates $K$ fold losses and typically has substantially smaller variance than a single hold-out evaluation. In this setting, the single-split bound remains valid but can be conservative. For practitioners who wish to approximate the uncertainty of the reported $K$-fold average more closely, we consider a simple CLT-style rescaling of the bound as a heuristic “practical” variance proxy, and evaluate its accuracy empirically. Across all tested noise regimes and hold-out sizes, this practical proxy closely tracks the Monte Carlo variance of the $K$-fold estimator well for $K\geq 4$ (Appendix Figure 9). While this heuristic does not carry the same finite-sample guarantee as the theoretical bound, it provides a useful operational approximation.
@@ A.5 Proof of Theorem 1
(Theorem 1 of Whittle (1960)) Let the random variables $z_{1},z_{2},\ldots,z_{n}$ be statistically independent, and each assume the values $\pm 1$ with respective probabilities $\frac{1}{2}$ . Then, for any set of real coefficients $b_{j}$, $s\geq 2$,